OpenStudy (anonymous):

2/9 divided by 1/4 Answer in lowest terms Answer today please!

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{2}{9} \div \frac{1}{4}= }\) \(\LARGE\color{black}{ \frac{2}{9} \times \frac{4}{1}= }\) \(\LARGE\color{black}{ \frac{2~~~\times~~~ 4 }{9~~~\times ~~~1} = }\) \(\LARGE\color{black}{ \frac{~~8~~~ }{~~9~~} . }\)

OpenStudy (anonymous):

^ *throws medal*

OpenStudy (solomonzelman):

no reason to....

OpenStudy (solomonzelman):

I certainly don't mind tho

OpenStudy (anonymous):

I like the way you explained. :)

OpenStudy (solomonzelman):

I can put up some more LateX, like \(\huge\color{ blue }{\huge {\bbox[5pt, lightyellow ,border:2px solid lightyellow ]{ ~~~~~~~~{\frak latex}✍ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ }}}\)

OpenStudy (solomonzelman):

oops, the symbols turned into a bunch of question marks ... hate that bug. The main thing is however, that you got the help needed, so bye:) See you....

OpenStudy (anonymous):

\(\LARGE\color{lime}{\ddot\smile}\) \(\LARGE\color{pink}{\ddot\smile}\) It would work perfect once you refresh your page. @SolomonZelman